OpenStudy (anonymous):

Find the first 5 terms of the sequence: a1 = 2, an = -3an-1

OpenStudy (anonymous):

okay this one is a bit trickier, but i think you can still do it!

OpenStudy (anonymous):

So, I though a2 would be 3 but, that didn't work. @robz8

OpenStudy (anonymous):

a2 = -3(a1)... so what is a1 again?

OpenStudy (anonymous):

Is it like this: \[a _{n} = -3a _{n-1}\]

OpenStudy (anonymous):

It says a1 is 2

OpenStudy (anonymous):

yes, so if a1 is 2... and a2 = -3 * 2..... what would that make a2?

OpenStudy (anonymous):

a2 would be -6? @robz8

OpenStudy (anonymous):

correct, so now that a2 = -6, the next part in the function is a3 = -3(a2) so what would a3 be?

OpenStudy (anonymous):

-27? @robz8

OpenStudy (anonymous):

nooooooo... not quite, try again. we know that a2 is -6, so a3 = -3*(-6)

OpenStudy (anonymous):

but is it written like how i wrote it? @Emilyh117 because the notation counts.

OpenStudy (anonymous):

Oooooo. It would be 18. @robz8

OpenStudy (anonymous):

there we go... now repeat that pattern 2 more times, and you have the first 5 terms of the recursive equation

OpenStudy (anonymous):

Okay. I started the next one and got -54. Then I got 162. @robz8

OpenStudy (anonymous):

that is correct! now you have the first 5 terms of the recursive equation!

OpenStudy (anonymous):

Yay! Thank you! :) @robz8

OpenStudy (anonymous):

yw! @Emilyh117